OpenStudy (anonymous):

Solve. 14x2 + 42x + 70 = 0

OpenStudy (anonymous):

factor out the 7 first

OpenStudy (anonymous):

divide them with 14

OpenStudy (anonymous):

14x^2/14 + 42x/14 + 70/14 = 0/14

OpenStudy (anonymous):

x^2 + 3x + 5 = 0 then try solving for x

OpenStudy (anonymous):

use quadratic formula \[x=[-b-+\sqrt{(b^2 -4ac)}/2a\], here you a=1,b=3,c=5 try them

OpenStudy (anonymous):

youl get an answer of x=-1.5+1.66i and x=-1.5-1.66i ans...

OpenStudy (anonymous):

remember that \[\sqrt{-1}=i\]

OpenStudy (anonymous):

let me know if you hve question

OpenStudy (anonymous):

im def. lost math is not my strongest subject. i just cant get it

OpenStudy (anonymous):

it just takes practise.....

OpenStudy (anonymous):

do this one til you get the answer

OpenStudy (anonymous):

you already know how to use the quadraic formula right?

OpenStudy (anonymous):

a little but not really

OpenStudy (anonymous):

ok try that till u get the answer